Solve the following problems using the correct format. Given first, what is asked for, and the formula together with the solution. 


  1. A gas occupies a volume of 400 cm3 at a pressure of 2.00 atm. What would be the new pressure if the volume is increased to 500 cm3 ?
  2. The volume of a given gas was 50 cm3 at 260C. After heat was applied to the gas, its volume increased to 80 cm3 . If the pressure was held constant, what is the new temperature of the gas? 
1
Expert's answer
2022-05-24T06:23:03-0400

1. A gas occupies a volume of 400 cm3 at a pressure of 2.00 atm. What would be the new pressure if the volume is increased to 500 cm3?


Given:

P1 = 2.00 atm

V1 = 400 cm3

T = constant

P2 = unknown

V2 = 500 cm3


Formula: P1V1 = P2V2


Solution:

Since the temperature and amount of gas remain unchanged, Boyle's law can be used.

Boyle's gas law can be expressed as: P1V1 = P2V2

To find the new pressure, solve the equation for P2:

P2 = P1V1 / V2

P2 = (2.00 atm × 400 cm3) / (500 cm3) = 1.60 atm

P2 = 1.60 atm


Answer: The new pressure will be 1.60 atm




2. The volume of a given gas was 50 cm3 at 26°C. After heat was applied to the gas, its volume increased to 80 cm3. If the pressure was held constant, what is the new temperature of the gas? 


Given:

V1 = 50 cm3

T1 = 26°C + 273.15 = 299.15 K

P = constant

V2 = 80 cm3

T2 = unknown


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the new temperature of the gas, solve the equation for T2:

T2 = V2T1 / V1

T2 = (80 cm3 × 299.15 K) / (50 cm3) = 478.64 K

T2 = 478.64 K (or 205.5°C)


Answer: The new temperature of the gas will be 478.64 K (or 205.5°C)